OpenStudy (anonymous):

Determine the solution set of (3x + 1)2 - 100 = 0. {3, -3} {3, -11/3} {-11/3, -3}

OpenStudy (anonymous):

substituting the {3,-11/3} the equation is satisfied for the both values individually. Also it can be satisfied by solving the equation 3x2+2x-33=0 and giving the roots as x=3 and x=-11/3

jhonyy9 (jhonyy9):

3x+1=+/-10 3x=+/-9 x=+/-3

OpenStudy (mertsj):

\[(3x+1)^2=100\] Take the square root of both sides

OpenStudy (mertsj):

Just like jhonyy9 says

jhonyy9 (jhonyy9):

thank you